4. Find the general solution of the differential equation dy/dx + y cot x = 1, recalling that cot x = cos x / sin x 


1
Expert's answer
2022-02-03T16:29:16-0500

We have;

"\displaystyle
\frac{dy}{dx}+y\cot (x)=1\Rightarrow \frac{dy}{dx}+y\frac{\cos(x)}{\sin(x)}=1"

The Integrating Factor (I.F) of the DE is;

"\displaystyle
e^{\int\frac{\cos (x)}{\sin(x)}\ dx}=e^{\ln|\sin(x)|}=\sin(x)"

Multiplying the I.F by the given DE yields;

"\displaystyle
\left(\frac{dy}{dx}+y\frac{\cos(x)}{\sin(x)}\right)\sin(x)=1\times\sin(x)\\
\Rightarrow \frac{d}{dx}(y\sin(x))=\sin(x)\\
\Rightarrow y\sin(x)=\int\sin(x)\ dx\\
\Rightarrow y\sin(x)=-\cos(x)+a,\text{ where a is an arbitrary constant.}\\
\Rightarrow y=\frac{a-\cos(x)}{\sin(x)}"